For parts a and b​, use technology to estimate the following. ​a) The critical value of t for a 98 ​% confidence interval with dfequals 9 . ​b) The critical value of t for a 90 ​% confidence interval with dfequals 111 . ​a) What is the critical value of t for a 98 ​% confidence interval with df equals 9​?
Answer #1
df confidence level t critical value excel function
a).9

98% = 0.98

alpha = (1-0.98) = 0.02

2.821 in any blank cell of excel type =T.INV.2T(0.02,9) press enter.
b).111

90% = 0.90

alpha = (1-0.90) = 0.10

1.659 in any blank cell of excel type =T.INV.2T(0.1,111) press enter.

***in case of doubt, comment below. And if u liked the solution, please like.